# Fin-RATE: A Real-world Financial Analytics and Tracking Evaluation Benchmark for LLMs on SEC Filings

### v1

@@ 4.1.2. Retrieval Settings
  • Dense Vector Retrieval: A semantic approach using general (all-MiniLM-L6-v2 (Wang et al., 2020)) and domain-specific (finance-embeddings-investopedia (FinLang, 2024)) embedding models to evaluate the impact of financial domain adaptation.
@@
  • Reranking Stage: To further refine the candidate set, we employ the bge-reranker-v2-m3 (Liu et al., 2025b; Chen et al., 2024) to re-scores the retrieved documents for high contextual relevance
@@ 4.2.1. Generation Evaluation Metrics
We evaluate model-generated answers using a unified LLM-as-Judge framework. Three LLMs with distinct architectures—GPT-5 (Singh et al., 2025), DeepSeek-V3.2 (DeepSeek-AI et al., 2025), and Qwen3-235B (Yang et al., 2025)—serve as independent judges. ach judge assesses the response strictly based on the provided reference answer or context, following standardized instructions that prohibit the use of external knowledge or speculative inference. To ensure reliable judgment, we derive weights through an alignment analysis (Appendix I) and fuse their outputs using weighted averaging for correctness labels and Likert scores, and a union rule for error-type attribution (Appendix J). The judge ensemble provides three assessments:
@@
Fine-grained Likert Scores assess each answer on five dimensions—information coverage, reasoning quality, factual consistency, clarity, and analytical depth with each rated on a 1–5 Likert scale.
Error Type Attribution captures the failure cause for all non-CORRECT responses, enabling diagnostic analysis. Definitions of error categories are provided in Appendix G.
@@ 4.3.1. Task-Level Challenge Landscape
Table 2, Figure 3, and Appendix F.2 demonstrate distinct challenges for EC-QA and LT-QA.
@@
LT-QA’s time mismatch (C3, 5.03%) and trend hallucination (B1-4, 8.43%) are the highest among all tasks. Intent misunderstanding (D1-1) further accounts for 35.38% of LT-QA errors, showing that models often treat trend-oriented questions as requests for yearly summaries. Besides, information coverage and reasoning scores decline by 0.88 and 1.03 compared with DR-QA. The model lacks mechanisms for temporal abstraction and cross-year integration. It processes yearly information as independent units rather than linking them into a continuous temporal sequence. Without a representation that captures how indicators evolve over time, the model cannot infer stable trends or causal relations and instead relies on isolated factual recall, which fragments its temporal reasoning. As a result, generated answers tend to include fabricated temporal claims, exhibit reduced clarity (-0.99), and become verbose without providing actionable trend insights.
@@ 4.3.2. Model category-Specific Behavior
Closed-source models exhibit the highest robustness across all tasks. Their Correct rate decreases only moderately from 39.71% in DR-QA to 33.33% in EC-QA, with Failure consistently below 1%. Radar plots show a nearly intact circular profile, with cross-dimensional degradation limited to 3.28%–12.49%. Such robustness likely stems from large parameter counts and extended training chains, along with engineering advances such as position encoding that support superior long-context parsing and minimize information loss. Despite low hallucination rates, closed-source models frequently misattribute reasoning steps. Hallucination (B1) errors decrease substantially from DR-QA to EC-QA, while Contradicts Evidence (B2) rises to 1,980 cases, revealing a core bottleneck in maintaining self-consistent reasoning, despite correct retrieval.
@@
Although achieving 27.43% Correct in DR-QA, their Failure rate spikes from 1.63% to 24.49% in EC-QA. EC-QA triggers simultaneous increases in multiple error types, with surges in Query Misunderstanding (D1), Hallucination (B1), and Contradicts Evidence (B2), along with a $6.8\times$ rise in Context Window Abuse (D2). This pattern arises because broad-domain pretraining and smaller capacity lead these models to rely on shallow lexical cues and short-range attention, preventing stable tracking of entities, temporal references, or section relations. With errors spreading across dimensions, the radar plots show uniform collapses across all evaluation axes as semantic drift and information loss cause early misalignments to compound along the reasoning chain.
@@
Correct rates drop from 23.92% in DR-QA to just 1.27% in EC-QA, with both Failure and Incorrect sharply rising. Radar plots reveal a lopsided profile, where nearly 50% of metrics degrade and Reasoning Chain drops by 49.16%. In DR-QA, Query Misunderstanding (D1) errors are only 804.5, comparable to top closed-source models. In EC-QA and LT-QA, it also produces fewer numerical errors (C1, C2, C4) than general open-source models. However, Hallucination (B1) rates remain consistently higher than in other model categories. Current finance-tuning pipelines rely on single-document supervision focused on lexical cues, terminology, and numerical patterns. While this enhances domain grounding, it also reinforces template-based guessing: models infer missing relations or numbers from familiar financial patterns, causing frequent hallucinations dominated by Entity Attribute Hallucination (B1-2).
@@ 4.3.3. Retrieval Behavior Across QA Tasks
Different QA scenarios show distinct retrieval preferences. DR-QA and LT-QA depend heavily on localized, explicitly anchored information, while EC-QA requires broader semantic correspondence across corporate narratives. BM25, which relies on lexical matching, achieves 40.4% recall in DR-QA and 94.3% accuracy in locating year-aligned evidence in LT-QA, outperforming VF at 30.4% and 74.7%. However, in EC-QA, VF retrieves multi-company evidence in 23.8% of cases, more than doubling BM25’s 10.4%. These patterns reflect fundamental differences in how the retrievers operate. BM25 prioritizes distinctive tokens such as years and financial terms, making it effective when explicit anchors exist. VF, in contrast, encodes semantic similarity and therefore better captures heterogeneous cross-company narratives, but loses precision on fine-grained temporal cues.
@@ 4.3.4. End-to-End RAG Failure Attribution
To evaluate end-to-end performance under retrieved context, we randomly select 500 examples for each QA type to evaluate model performance under retrieved context. For EC-QA and LT-QA, we restrict our evaluation to the subset of questions that contain at most five gold chunks.
@@
To further analyze task-specific failure patterns under RAG, we categorize retrieval-stage errors into three types: Missing Evidence (no gold chunk retrieved), Sorting Failure (gold retrieved but ranked below top-3), and Distractor Evidence (gold appears within the top-3 but accounts for no more than 20% of the top-5 context). EC-QA exhibits the most severe breakdowns. Figure 5 shows that Missing Evidence reaches 75.44%. Moreover, 57.6% of multi-company questions fail to retrieve all relevant entities, leaving key company evidence absent. This forces the model to conduct comparative reasoning over severely asymmetric or biased context. LT-QA shows a distinct error profile. The 7.64% Sorting Failure and 4.96% Distractor Evidence reflect frequent temporal misalignment, with correct-year gold chunks pushed down by similar paragraphs from other years or entangled with multi-year narratives in the top-5 results.
@@ 4.3.5. Hierarchical Retrieval Improvements
We evaluate hierarchical retrieval using R@10, P@1, and MRR, to measure overall coverage, top-ranked accuracy, and ranking quality. Hierarchical retrieval substantially improves both evidence coverage and ranking quality by constraining search within entity–year buckets. Compared to BM25+Reranker, the hierarchical retriever improves P@1 by 14.6 points in DR-QA, Recall@10 by 4.29 points in LT-QA, and both Recall@10 and MRR by 12.76 and 16.64 points respectively in EC-QA. Entity-aware routing also offers clear benefits: company hit rate in EC-QA rises from 13.2% to 52.9%, year coverage in LT-QA increases from 94.6% to 98.7%. These improvements suggest that future RAG systems should adopt more structure-aware, entity- and time-guided retrieval.
@@ 5. Conclusion
In this work, we introduce Fin-RATE, a benchmark constructed from real-world SEC filings to evaluate LLMs’ financial reasoning across detail interpretation, enterprise comparison, and longitudinal tracking. Fin-RATE highlights persistent challenges in aligning time-sensitive evidence, synthesizing multi-entity disclosures, and performing longitudinal financial reasoning across fiscal years. While LLMs perform well when reasoning over isolated passages, performance deteriorates sharply in tasks requiring interlinked evidence, such as cross-entity comparison or multi-year trend analysis. In EC-QA and LT-QA, we observe notable drops in clarity and factual consistency, driven by hallucinated comparisons, trend misalignment, and intent misunderstanding. Among different model families, finance-tuned models emerge as the most sensitive to task complexity. While they demonstrate strong command of domain-specific terminology and numerical accuracy, they are prone to hallucinations during complex multi-entity reasoning. This fragility underscores the importance of integrating long-form contexts and deeper logical dependencies into the fine-tuning process.
Appendix A LLM Focus Guidance for Detail & Reasoning QA
Appendix B LLM Focus Guidance for Enterprise Comparison QA
Appendix C LLM Focus Guidance for Longitudinal Tracking QA
Appendix D Detailed Corpus and Task Statistics
@@
D.1. Document Length and Type Distribution
@@ 60 Services 5 Construction, Real Estate, Hospitality
D.2. Document Type Distribution by Question Category
@@
D.3. Question Type and Context Analysis
@@
Appendix E Illustrative Example Questions
E.1. Detail & Reasoning QA
Figure 9. Example for Detail & Reasoning QA (DR) question
E.2. Enterprise Comparison QA
Figure 10. Example for Enterprise Comparison (EC) question
E.3. Longitudinal Tracking QA
Figure 11. Example for Longitudinal Tracking (LT) question
Appendix F Supplementary Evaluation Results
F.1. Full-Label Evaluation under Retrieved Context
@@
F.2. Likert-Scale Evaluation across Five Dimensions
@@
F.3. Error Pattern Analysis across Model Families
@@
Appendix G Error Analysis
G.1. Retrieval-Related Errors
G.2. Generation-Related Errors
G.3. Finance-Specific Numerical & Semantic Errors
G.4. Query and Contextual Understanding Errors
Appendix H Human Verification Criteria
H.1. Factual Alignment
H.2. Contribution of Referenced Chunks
H.3. Numerical Verification
Appendix I Consistency Verification Between LLM and Human Evaluation
Judge Model DR-QA EC-QA LT-QA
DeepSeek-V3.2 84.29 82.99 81.35
Qwen3-235B 84.10 84.34 84.67
GPT-5 79.57 90.22 90.07
Multi-Judge (Weighted) 88.82 90.41 90.46
Appendix J Fusion Strategy for Multi-Judge Evaluation
This section details the fusion strategies applied to the evaluation outputs of multiple LLM Judges, including correctness labels, Likert scores, and error type attributions. All fusion methods are based on the optimal task-specific weights derived from the human alignment experiment described in Appendix I.
For each QA instance, we follow the same procedure as in the grid search setting of the alignment experiment. Discrete correctness labels are first mapped to ordinal numeric scores, where Failure is mapped to 0, Incorrect to 1, Partial to 3, and Correct to 5. Based on the task-specific weights optimized in Appendix I, we then compute a weighted average across the three Judges’ scores to obtain a continuous fused score. This score is mapped back to a final discrete label using the following rule: a score of 0 yields Failure, a score of 1 yields Incorrect, scores of 2–3 yield Partial, and scores of 4–5 yield Correct.
@@ Appendix I Consistency Verification Between LLM and Human Evaluation
We evaluate a diverse set of large language models (LLMs) covering different training paradigms, parameter scales, and domain specializations, as summarized in Table 22. Open-source models are evaluated either via local inference on NVIDIA A100/A6000 GPUs ($\leq 70$B) or through APIs for larger models (e.g., DeepSeek and Qwen3-235B), and all models use a unified decoding temperature of 0.7.
@@
  • Requirements for question uniqueness and document-specificity

### v2

@@ 4.1.2. Retrieval Settings
  • Dense Vector Retrieval: A semantic approach using the most downloaded general (all-MiniLM-L6-v2 (Wang et al., 2020)) and domain-specific (finance-embeddings-investopedia (FinLang, 2024)) embedding models on HuggingFace to evaluate financial domain adaptation’s impact.
@@
  • Reranking Stage: To further refine the candidate set, we employ the bge-reranker-v2-m3 (Liu et al., 2025; Chen et al., 2024) to re-scores the retrieved documents for high contextual relevance
@@ 4.2.1. Generation Evaluation Metrics
We evaluate model-generated answers using a unified LLM-as-Judge framework. Three LLMs with distinct architectures—GPT-5 (Singh et al., 2025), DeepSeek-V3.2 (DeepSeek-AI et al., 2025), and Qwen3-235B (Yang et al., 2025)—serve as independent judges. Each judge assesses the response strictly based on the provided reference answer or context, following standardized instructions that prohibit the use of external knowledge or speculative inference. To ensure reliable judgment, we derive weights through an alignment analysis (Appendix F) and fuse their outputs using weighted averaging for correctness labels and Likert scores, and a union rule for error-type attribution (Appendix G). The judge ensemble provides three assessments:
@@
Fine-grained Likert Scores assess each answer on five dimensions—information coverage, reasoning quality, factual consistency, clarity, and analytical depth, with each rated on a 1–5 Likert scale.
Error Type Attribution captures the failure cause for all non-CORRECT responses, enabling diagnostic analysis. Definitions of error categories are provided in Appendix H.
@@ 4.3.1. Task-Level Challenge Landscape
Table 2, Figure 3, and Appendix I.2 demonstrate distinct challenges for EC-QA and LT-QA.
@@
LT-QA’s time mismatch (C3, 5.03%) and trend hallucination (B1-4, 8.43%) are the highest among all tasks. Intent misunderstanding (D1-1) further accounts for 35.38% of LT-QA errors, showing that models often treat trend-oriented questions as requests for yearly summaries. Besides, information coverage and reasoning scores decline by 0.88 and 1.03, respectively, compared with DR-QA. The model lacks mechanisms for temporal abstraction and cross-year integration. It processes yearly information as independent units rather than linking them into a continuous temporal sequence. Without a representation that captures how indicators evolve over time, the model cannot infer stable trends or causal relations and instead relies on isolated factual recall, which fragments its temporal reasoning. As a result, generated answers tend to include fabricated temporal claims, exhibit reduced clarity (-0.99), and become verbose without providing actionable trend insights.
@@ 4.3.2. Model category-Specific Behavior
Closed-source models exhibit the highest robustness across all tasks. Their Correct rate decreases only moderately from 39.69% in DR-QA to 33.28% in EC-QA, with Failure consistently below 1%. Radar plots show a nearly intact circular profile, with cross-dimensional degradation limited to 3.28%–12.49%. Such robustness likely stems from large parameter counts and extended training chains, along with engineering advances such as position encoding that support superior long-context parsing and minimize information loss. Despite low hallucination rates, closed-source models frequently misattribute reasoning steps. Hallucination (B1) errors decrease substantially from DR-QA to EC-QA, while Contradicts Evidence (B2) rises to 1,980 cases, revealing a core bottleneck in maintaining self-consistent reasoning, despite correct retrieval.
@@
Although achieving 27.42% Correct in DR-QA, their Failure rate spikes from 1.63% to 24.49% in EC-QA. EC-QA triggers simultaneous increases in multiple error types, with surges in Query Misunderstanding (D1), Hallucination (B1), and Contradicts Evidence (B2), along with a $6.8\times$ rise in Context Window Abuse (D2). This pattern arises because broad-domain pretraining and smaller capacity lead these models to rely on shallow lexical cues and short-range attention, preventing stable tracking of entities, temporal references, or section relations. With errors spreading across dimensions, the radar plots show uniform collapses across all evaluation axes as semantic drift and information loss cause early misalignments to compound along the reasoning chain.
@@
Correct rates drop from 23.91% in DR-QA to just 1.27% in EC-QA, with both Failure and Incorrect sharply rising. Radar plots reveal a lopsided profile, where nearly 50% of metrics degrade, and Reasoning Chain drops by 49.16%. In DR-QA, Query Misunderstanding (D1) errors are only 804.5, comparable to top closed-source models. In EC-QA and LT-QA, it also produces fewer numerical errors (C1, C2, C4) than general open-source models. However, Hallucination (B1) rates remain consistently higher than in other model categories. Current finance-tuning pipelines rely on single-document supervision focused on lexical cues, terminology, and numerical patterns. While this enhances domain grounding, it also reinforces template-based guessing: models infer missing relations or numbers from familiar financial patterns, causing frequent hallucinations dominated by Entity Attribute Hallucination (B1-2).
@@ 4.3.3. Retrieval Behavior Across QA Tasks
Different QA scenarios show distinct retrieval preferences. DR-QA and LT-QA depend heavily on localized, explicitly anchored information, while EC-QA requires broader semantic correspondence across corporate narratives. BM25, which relies on lexical matching, achieves 41.2% recall in DR-QA and 94.3% accuracy in locating year-aligned evidence in LT-QA, outperforming VF at 27.4% and 74.7%. However, in EC-QA, VF retrieves multi-company evidence in 23.8% of cases, more than doubling BM25’s 10.4%. These patterns reflect fundamental differences in how the retrievers operate. BM25 prioritizes distinctive tokens such as years and financial terms, making it effective when explicit anchors exist. VF, in contrast, encodes semantic similarity and therefore better captures heterogeneous cross-company narratives, but loses precision on fine-grained temporal cues.
@@ 4.3.4. End-to-End RAG Failure Attribution
To evaluate end-to-end performance under retrieved context, we randomly select 500 examples for each QA type to evaluate model performance under retrieved context. For EC-QA and LT-QA, we restrict our evaluation to the subset of questions that contain at most five gold chunks, which serve as the ground-truth context.
@@
To further analyze task-specific failure patterns under RAG, we categorize retrieval-stage errors into three types: Missing Evidence (no gold chunk retrieved), Sorting Failure (gold retrieved but ranked below top-3), and Distractor Evidence (gold appears within the top-3 but accounts for no more than 20% of the top-5 context). EC-QA exhibits the most severe breakdowns. Figure 5 shows that Missing Evidence reaches 75.44%. Moreover, 57.6% of multi-company questions fail to retrieve all relevant entities, leaving key company evidence absent. This forces the model to conduct comparative reasoning over severely asymmetric or biased contexts. LT-QA shows a distinct error profile. The 7.64% Sorting Failure and 4.96% Distractor Evidence reflect frequent temporal misalignment, with correct-year gold chunks pushed down by similar paragraphs from other years or entangled with multi-year narratives in the top-5 results.
@@ 4.3.5. Hierarchical Retrieval Improvements
We evaluate hierarchical retrieval using R@10, P@1, and MRR, to measure overall coverage, top-ranked accuracy, and ranking quality. Hierarchical retrieval substantially improves both evidence coverage and ranking quality by constraining search within entity–year buckets. Compared to BM25+Reranker, the hierarchical retriever improves P@1 by 5.64 points in DR-QA, R@10 by 4.29 points in LT-QA, and both R@10 and MRR by 12.76 and 16.64 points respectively in EC-QA. Entity-aware routing also offers clear benefits: company hit rate in EC-QA rises from 13.2% to 52.9%, year coverage in LT-QA increases from 94.6% to 98.7%. These improvements suggest that future RAG systems should adopt more structure-aware, entity- and time-guided retrieval.
@@ 5. Conclusion
In this work, we introduce Fin-RATE, a benchmark constructed from real-world SEC filings to evaluate LLMs’ financial reasoning across detailed interpretation, enterprise comparison, and longitudinal tracking. Fin-RATE highlights persistent challenges in aligning time-sensitive evidence, synthesizing multi-entity disclosures, and performing longitudinal financial reasoning across fiscal years. While LLMs perform well when reasoning over isolated passages, performance deteriorates sharply in tasks requiring interlinked evidence, such as cross-entity comparison or multi-year trend analysis. In EC-QA and LT-QA, we observe notable drops in clarity and factual consistency, driven by hallucinated comparisons, trend misalignment, and intent misunderstanding. Among different model families, finance-tuned models emerge as the most sensitive to task complexity. While they demonstrate strong command of domain-specific terminology and numerical accuracy, they are prone to hallucinations during complex multi-entity reasoning. This fragility underscores the importance of integrating long-form contexts and deeper logical dependencies into the fine-tuning process.
Appendix A Detailed Corpus and Task Statistics
@@
A.1. Document Length and Type Distribution
@@ 60 Services 5 Construction, Real Estate, Hospitality
A.2. Document Type Distribution by Question Category
@@
A.3. Question Type and Context Analysis
@@
Appendix B LLM Focus Guidance for Detail & Reasoning QA
Appendix C LLM Focus Guidance for Enterprise Comparison QA
Appendix D LLM Focus Guidance for Longitudinal Tracking QA
Appendix E Human Verification Criteria
E.1. Factual Alignment
E.2. Contribution of Referenced Chunks
E.3. Numerical Verification
Appendix F Consistency Verification Between LLM and Human Evaluation
Judge Model DR-QA EC-QA LT-QA
DeepSeek-V3.2 84.29 82.99 81.35
Qwen3-235B 84.10 84.34 84.67
GPT-5 79.57 90.22 90.07
Multi-Judge (Weighted) 88.82 90.41 90.46
Appendix G Fusion Strategy for Multi-Judge Evaluation
This section details the fusion strategies applied to the evaluation outputs of multiple LLM Judges, including correctness labels, Likert scores, and error type attributions. All fusion methods are based on the optimal task-specific weights derived from the human alignment experiment described in Appendix F.
For each QA instance, we follow the same procedure as in the grid search setting of the alignment experiment. Discrete correctness labels are first mapped to ordinal numeric scores, where Failure is mapped to 0, Incorrect to 1, Partial to 3, and Correct to 5. Based on the task-specific weights optimized in Appendix F, we then compute a weighted average across the three Judges’ scores to obtain a continuous fused score. This score is mapped back to a final discrete label using the following rule: a score of 0 yields Failure, a score of 1 yields Incorrect, scores of 2–3 yield Partial, and scores of 4–5 yield Correct.
Appendix H Error Type Taxonomy
H.1. Retrieval-Related Errors
H.2. Generation-Related Errors
H.3. Finance-Specific Numerical & Semantic Errors
H.4. Query and Contextual Understanding Errors
Appendix I Supplementary Evaluation Results
I.1. Full-Label Evaluation under Retrieved Context
@@
I.2. Likert-Scale Evaluation across Five Dimensions
@@
I.3. Error Pattern Analysis across Model Families
@@
Appendix J Illustrative Example Questions
J.1. Detail & Reasoning QA
Figure 12. Example for Detail & Reasoning QA (DR) question
J.2. Enterprise Comparison QA
Figure 13. Example for Enterprise Comparison (EC) question
J.3. Longitudinal Tracking QA
Figure 14. Example for Longitudinal Tracking (LT) question
@@
We evaluate a diverse set of large language models (LLMs) covering different training paradigms, parameter scales, and domain specializations, as summarized in Table 22. Open-source models are evaluated either via local inference on NVIDIA A100/A6000 GPUs ($\leq 70$B) or through APIs for larger models like Qwen3-235B, and all models use a unified decoding temperature of 0.7.
@@
  • Requirements for question uniqueness and document relevance